5. What is the hydroxide ion concentration is a solution with pH = 8.55?
Answer #1

pH = 8:55 pon--10g Cor] 1000M = con] - pH + pot = 14 son – 14 - 314-8.55 : 1pon = 5.45] [omi] = 10-pon 210-5.45 [on] = 3.5488